40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 27 23:49:56 2014, Koji, Update, General, IR beam found, TTs not aligned well yet 
    Reply  Mon Jun 30 16:55:24 2014, ericq, Update, General, IFO (mostly) aligned 
       Reply  Mon Jun 30 22:06:55 2014, Jenne, Update, ASC, IFO (mostly) aligned - ASS working 
          Reply  Wed Jul 2 16:18:32 2014, Jenne, Update, ASC, IFO aligned, PRMI + 2 arms achieved 
Message ID: 10113     Entry time: Mon Jun 30 16:55:24 2014     In reply to: 10110     Reply to this: 10114
Author: ericq 
Type: Update 
Category: General 
Subject: IFO (mostly) aligned 

IR Alignment of the IFO is recovered! Green alignment could use some attention. 

Arms and PRC hold well, powers are within normal ranges. (TR[X,Y] >.9, POP110I > 400)

Details:

  • Starting from where Koji had brought things (i.e. beams on REFL, AS), I gradually stepped the PRM back to where I had a PRC lock right before the bootfest, using SUSPIT_IN and SUSYAW IN, while stepping the TT pointing (arbitrarily shifting TT1 and TT2) to keep the REFL camera spot centered. 
  • Green was locked fairly well to the Yarm (GTRY = 0.8), so I knew it supported a mode. I misaligned ITMX to keep the Xarm out of the way. 
  • Adjusting TT1 and TT2, I was able to get some light on the POP camera, while keeping REFL spot visible with small adjustments to PRM
  • I set up the PRM to set up some PRM-ITMY retroreflection, and saw tiny flashes in TRY. 
  • From here, I wandered around with the tip-tilts to try and get better arm pointing, while tweaking PRM for the REFL spot and retroflection, and BS for the AS spot. I got this to TRY flashes of ~0.3

Then, Jenne took over, worked some alignment magic, and did the hard part of getting the Yarm locked and ASS'd. 

  • I then took back over and got the Xarm locked and ASS'd.
  • I saved relevant mirror positions, and set up the PRMI, got it locked, saved PRM position. 
  • Jiggled the BS/PRM oplev HeNe power connector, it turns on now, oplevs are working. 
  • Similar to TT1, TT2 gains for PIT and YAW are now 300. Strictly speaking, they don't have to be, but PIT would be very nearly railed, and I changed YAW so that all four gains would be consistent.

Caveats:

Green no longer locks to 00 on the Y-arm, and the X-arm green transmission isn't the best despite PZT fiddling (~.6). Also, when green is locked to the Xarm, I see a distinct circular spot on the GTRY camera, with Y green and PSL green shutters closed. 

While Jenne used Yarm ASS successfully, when I run it now, it slowly pushes things out of alignment, and two of the traces (YARM_ETM_YAW_L_DEMOD_I_OUTPUT, and the same for ITM) have a reasonable constant offset that doesn't move away. 

 

ELOG V3.1.3-